# A Real-Time Adaptive Multi-Stream GPU System for Online Approximate Nearest Neighborhood Search

### v1

@@ 1. Introduction
In this article, we introduce a novel Real-Time Adaptive Multi-Stream System for Online GPU Approximate Nearest Neighbor Search (RTMS-GANNS). Our innovation includes a well-designed dynamic vector operation algorithm along with a novel execution mode which leverages the parallelism inherent in GPU streams, enabling a high-performance concurrent execution of vector search and real-time vector insertion. Our contributions can be summarized as follows:
@@ 3.1. Memory Block Based Real-time Vector Insertion Algorithm
For each memory block $m$, it may contain $|ml|$ vectors, which of limit $T_{m}$ can be pre-set by users. The structure of a memory block is also illustrated in Figure 1b. It consists of a header $mh$, core id list $ml$, and core vectors $mv$. The header $mh$ contains important information that indicates the address of the prev_header, next_header and block_info (vector capacity, vector size, etc.). The core id list $ml$ saves the vector ids belong to the memory block of this cluster. The core vectors $mv$ in each memory block are arranged interleaved by 32 dimensions, similar to Faiss/Raft, determined by the thread number in a GPU warp to enable collapsed read instructions.
@@
The rearrangement algorithm is illustrated in algorithm.3. The core idea of rearrangement is to merge two split memory block together so that the vectors are continuous and a header jump is eliminated. In a certain memory block list $m^{\prime}$, we exchange the memory block next to the first memory block with linked memory block using a extra temp memory segment and then update the first block header. It may happen that the memory block we want to exchange is a merged memory block, under certain circumstance, we should first split it and do merge on it recursively.
@@ 3.3. Deployment Detail
The offline segment is dedicated to processing prepared data, while the online segment orchestrates real-time vector insertion operations. Managed by multiple CPU threads and interfacing with Kafka for message ingestion, the online segment implements a dynamic batching strategy, aggregating vectors either every second or upon reaching a threshold of 128 insertions, with a cap at 1024, before dispatching the aggregated batch to the GPU.
@@ 4. Experiment
  • RTMS-GANNS(Proposed Method): The first dynamic vector insertion system for ivfflat and ivfpq on the GPU, supporting multi-stream parallelism with retrieval.
@@ 4.1. Performance on different QPS
We can first conclude that the RTMS-GANNS (proposed method) not only achieves latency benefits on small datasets (Figure.3abc) but also shows significant improvements on large datasets (Figure.3def).
Firstly, RTMS-GANNS can outperform previous systems with up to a 40% reduction in latency even at low $\text{QPS}_{insertion}$, benefiting from dynamic vector insertion algorithm and multi-stream parallelism. Requests are immediately processed upon arrival without waiting for the previous task to complete, reducing latency by at least 1ms in Figure.3a and d. Secondly, as QPS increases, RTMS-GANNS has the lowest growth rate in latency, with a maximum reduction of over 80% (Figure.3ef). Except for a slight increase in Figure.3f due to the higher computational load of large datasets, it shows nearly linear growth under other parameters. Moreover, other systems timeout a lot (reflecting the $\text{latency}_{\text{avg}}$ is over 20ms) when $\text{QPS}_{search}$ is over 5000 and $\text{QPS}_{insertion}$ is over 1000, demonstrating that RTMS-GANNS’s dynamic vector insertion algorithm can efficiently run within the multi-stream parallel framework without causing any delay.
Among the four systems, the CPU-based system has the highest latency due to its inferior performance compared to the GPU. However, at lower $\text{QPS}_{search}$ on DSSMRT40M, it exhibits a linear increase in latency, which makes it outperform Faiss at $\text{QPS}_{insertion}$ = 2000 in Figure.3d. This observation aligns with our practical experience, where real-time insertions on CPU systems do not significantly affect normal retrieval performance. Additionally, it was found that both the retrieval and insertion performance of Faiss lag behind Raft and RTMS-GANNS. This is because the add operation in Faiss requires copying the vector list to the CPU and then back to the GPU, which is highly inefficient for large datasets, and the search operation is slower.

### v2

@@ 1. Introduction
In this article, we introduce a novel Real-Time Adaptive Multi-Stream System for Online GPU Approximate Nearest Neighbor Search (RTAMS-GANNS). Our innovation includes a well-designed dynamic vector operation algorithm along with a novel execution mode which leverages the parallelism inherent in GPU streams, enabling a high-performance concurrent execution of vector search and real-time vector insertion. Our contributions can be summarized as follows:
@@ 3.1. Memory Block Based Real-time Vector Insertion Algorithm
For each memory block $m$, it may contain $|ml|$ vectors, which of limit $T_{m}$ can be pre-set by users. The structure of a memory block is also illustrated in Figure 1b. It consists of a header $mh$, core id list $ml$, and core vectors $mv$. The header $mh$ contains important information that indicates the address of the prev_header, next_header and block_info (vector capacity, vector size, etc.). The core id list $ml$ saves the vector ids belong to the memory block of this cluster. The core vectors $mv$ in each memory block are arranged interleaved by 32 dimensions, similar to Faiss/Raft, determined by the thread number in a GPU warp to enable coalesced read instructions.
@@
The rearrangement algorithm is illustrated in algorithm.3. The core idea of rearrangement is to merge two split memory block together so that the vectors are continuous and a header jump is eliminated. In a certain memory block list $m^{\prime}$, we exchange the memory block next to the first memory block with linked memory block using a extra temp memory segment and then update the first block header. It may happen that the memory block we want to exchange is a merged memory block, under certain circumstance, we should first split it and do merge on it recursively or iteratively.
@@ 3.3. Deployment Detail
The offline segment is dedicated to processing prepared data, while the online segment orchestrates real-time vector insertion operations. Managed by multiple CPU threads and interfacing with Kafka for message ingestion, the online segment implements a dynamic batching strategy, aggregating vectors either every second or upon reaching a threshold at multiples of 128 insertions, with a cap at 1024, before dispatching the aggregated batch to the GPU.
@@ 4. Experiment
  • RTAMS-GANNS(Proposed Method): The first dynamic vector insertion system for ivfflat and ivfpq on the GPU, supporting multi-stream parallelism with retrieval.
@@ 4.1. Performance on different QPS
We can first conclude that the RTAMS-GANNS (proposed method) not only achieves latency benefits on small datasets (Figure.3abc) but also shows significant improvements on large datasets (Figure.3def).
Firstly, RTAMS-GANNS can outperform previous systems with up to a 40% reduction in latency even at low $\text{QPS}_{insertion}$, benefiting from dynamic vector insertion algorithm and multi-stream parallelism. Requests are immediately processed upon arrival without waiting for the previous task to complete, reducing latency by at least 1ms in Figure.3a and d. Secondly, as QPS increases, RTAMS-GANNS has the lowest growth rate in latency, with a maximum reduction of over 80% (Figure.3ef). Except for a slight increase in Figure.3f due to the higher computational load of large datasets, it shows nearly linear growth under other parameters. Moreover, other systems timeout a lot (reflecting the $\text{latency}_{\text{avg}}$ is over 20ms) when $\text{QPS}_{search}$ is over 5000 and $\text{QPS}_{insertion}$ is over 1000, demonstrating that RTAMS-GANNS’s dynamic vector insertion algorithm can efficiently run within the multi-stream parallel framework without causing any delay.
Among the four systems, the CPU-based system has the highest latency due to its inferior performance compared to the GPU. However, at lower $\text{QPS}_{search}$ on DSSMRT40M, it exhibits a linear increase in latency, which makes it outperform Faiss at $\text{QPS}_{insertion}$ = 2000 in Figure.3d. This observation aligns with our practical experience, where real-time insertions on CPU systems do not significantly affect normal retrieval performance. Additionally, it was found that both the retrieval and insertion performance of Faiss lag behind Raft and RTAMS-GANNS. This is because the add operation in Faiss requires copying the vector list to the CPU and then back to the GPU, which is highly inefficient for large datasets, and the search operation is slower.